OpenStudy (anonymous):

Find the inverse of the function. f(x) = 6x3 + 3

OpenStudy (anonymous):

\[f^1=\frac{ \sqrt[3]{36(x-3)} }{ 6} \]

OpenStudy (anonymous):

f(x) = 6x^3 + 3 let f(x) = y y = 6x^3 + 3 solve the function for x. then replace all the x's with y's and all the y's with x's.